the product of two positive numbers is 5/6 and their quotient is 15/8. calculate the numbers
Answer by Aldorozos(172)

x*y = 5/6
x/y = 15/8
In x/y = 15/8 multiply both sides by y to get ride of the denominator and find out what x is in terms of y
x = 15/8 * y
Now that we know what x is in terms of Y we replace x in x*y = 5/6 and we get
(15/8 * y)* y = 5/6
y^2 *15/8 = 5/6 Therefore y^2 = 5/6 *8/15 (here we multiplied both sides by 15/8 to isolate y^2)
Now we have y^2 = 4/9 Therefore y = +2/3 and -2/3
Now that we have y we replace y in x*y = 5/6 with +2/3 and -2/3 and we get two values for x which are 5/4 and -5/4